Let X denotes the number of hours a person uses a mobile and the probability distribution of X is as
|
X |
0 |
1 |
2 |
3 |
4 |
|
P(X) |
0.1 |
K |
2K |
2K |
K |
Then the value of K is
Answer & explanation
Correct answer: option 1
The correct answer is Option (1) → 0.15
Given the probability distribution:
| $X$ | 0 | 1 | 2 | 3 | 4 |
|---|---|---|---|---|---|
| $P(X)$ | 0.1 | $K$ | $2K$ | $2K$ | $K$ |
Sum of all probabilities must equal 1:
$0.1 + K + 2K + 2K + K = 1$
$0.1 + 6K = 1$
$6K = 1 - 0.1 = 0.9$
$K = \frac{0.9}{6} = 0.15$